### v1

# DiffImp: Efficient Diffusion Model for Probabilistic Time Series Imputation with Bidirectional Mamba Backbone

Abstract
Probabilistic time series imputation has been widely applied in real-world scenarios due to its ability to estimate uncertainty of imputation results. Meanwhile, denoising diffusion probabilistic models (DDPMs) have achieved great success in probabilistic time series imputation tasks with its power to model complex distributions. However, current DDPM-based probabilistic time series imputation methodologies are confronted with two types of challenges: 1)  The backbone modules of the denoising parts are not capable of achieving sequence modeling with low time complexity. 2) The architecture of denoising modules can not handle the inter-variable and bidirectional dependencies in the time series imputation problem effectively. To address the first challenge, we integrate the computational efficient state space model, namely Mamba, as the backbone denosing module for DDPMs. To tackle the second challenge, we carefully devise several SSM-based blocks for bidirectional modeling and inter-variable relation understanding. Experimental results demonstrate that our approach can achieve state-of-the-art time series imputation results on multiple datasets, different missing scenarios and missing ratios.
1 Introduction
The analysis of time series can model the intrinsic patterns within time-series data, thus providing robust support for decision-making in various fields, such as meteorology McGovern et al. (2011); Karevan & Suykens (2020), financial analysis Xiang et al. (2022); Owusu et al. (2023); Bai et al. (2020), healthcare Morid et al. (2023); Poyraz & Marttinen (2023) and power systems Tzelepi et al. (2023); Zhou et al. (2021). To enhance the reliability of analytical outcomes, it is critical to ensure the integrity of time series. However, due to various reasons such as device failures, human errors, and privacy protection, time series data can easily be incomplete with missing observations at different timestamps.
Time series imputation methods aim to estimate the values of missing points based on the observed points in incomplete time series, thereby restoring the integrity of the time series while preserving its original statistical properties. According to the ability to provide uncertainty of estimations, time series imputation methods can be categorized into the following two perspectives: 1) Deterministic Cao et al. (2018); Cini et al. (2022); Du et al. (2023), and 2)  Probabilistic Chen et al. (2023); Kim et al. (2023); Luo et al. (2018) imputation methods. Probabilistic time series imputation is particularly important in dealing with complex and uncertain data environments, as it provides a quantification of uncertainty for the imputations. The key to probabilistic imputation lies in modeling the posterior distribution. Existing probabilistic time series imputation methods include Gaussian Process and Variational Autoencoder-based methods Fortuin et al. (2020), Normalization Flow-based methods Rasul et al. (2021), and Diffusion-based methods Tashiro et al. (2021). Among these, the Diffusion-based method has emerged as the optimal choice for probabilistic time series due to their accuracy in posterior modeling and adaptability to different scenarios and various types of time series data.
When selecting a denoising backbone in the diffusion model, the following two key factors need to be considered: 1) Model compatibility, and 2) Time complexity. Specifically, the missing observations in time series have correlations with their neighbors on both sides, so it is crucial to design a model by considering information from neighbors of both sides. Moreover, it is also essential to accurately capture the properties of time series, such as global dependencies and channel correlations. Three mainstream denoising backbones are widely used in diffusion models for time series imputation: 1) Convoluational Neural Networks (CNNs)-, 2) Transformer- and 3) State-Space Model (SSM)-based backbones. Given a time series with a length of $L$, the CNNs-based backbone can capture partial information from the neighbors within the receptive fields and has $\mathcal{O}(L)$ time complexity. The transformer-based backbone can model temporal dependencies across the entire time series but is with quadratic time complexity $\mathcal{O}(L^{2})$. The SSM backbone has a linear time complexity, $\mathcal{O}(L)$, but it falls short in capturing the information from one side of the neighbor. Moreover, all these backbones fail to capture the channel dependencies in time series. The comparison results of existing backbones and our method in terms of various dependencies and time complexity are presented in Table.1.
Table 1: Comparison of our method and existing methods in modeling dependencies and time complexity. The results show that our method achieves the most comprehensive data modeling with the lowest time complexity.
Backbone Model Global Dependency Time Complexity Channel Dependency Inter-sequence Dependency
CNN Local $\mathcal{O}(L)$ Independent Unidirectional
Transformer Global $\mathcal{O}(L^{2})$ Independent Unidirectional
SSM Partial $\mathcal{O}(L)$ Independent Unidirectional
DiffImp (Ours) Global $\mathcal{O}(L)$ Dependent Bidirectional
In this paper, we propose an efficient diffusion-based framework for probabilistic time series imputation to address the drawbacks in existing backbones of time series imputation, we name it DiffImp. To ensure linear complexity, we choose the SSM-based model as the backbone of our framework, which is Mamba Dao & Gu (2024) to be more specific. To enable Mamba to capture information from both sides of the missing values, we then propose a Bidirectional Attention Mamba block (BAM) that is more applicable to time series imputation task. To incorporate bidirectional dependencies, we design a learnable weight module inside the BAM block. This module learns the weights of all points within the sequence, facilitating the modeling of dependencies at different distances. Next, we propose a Channel Mamba Block (CMB) to capture the dependencies among different channels in a time series. Specifically, we treat the variables across different channels in the time series as a sequence of variables and employ the Mamba model alongside the channel dimension, so inter-dependencies among channels can be modeled.
Our contributions are summarized as follows:
  • We propose DiffImp, an efficient diffusion-based model for the time series imputation task. We leverage the advantages of the bidirectional structure to devise blocks that can model sequence information and channel dependencies as the denoising backbone for the diffusion modules for performance improvements.
  • We dive deeply into the application of Mamba blocks by exploring the application of Mamba as the backbone block of the DDPMs for probabilistic time series imputation. We integrate the Mamba-based blocks to the diffusion models, therefore achieving effectively time series modeling with linear time complexity.
  • We conduct experiments on multiple real-world datasets for both time series imputation and time series forecasting tasks. The experimental results demonstrate that our approach achieves state-of-the-art performance across several datasets, different missing scenarios and missing ratios.
2 Preliminaries
2.1 State Space Models
State Space Models (SSMs) are an emerging approach to model sequential data, which is implemented by finding out state representations to model the relationship between input and output sequences. A SSM receives a one-dimensional sequence $X\in\mathbb{R}^{L}$ as the input and outputs a corresponding sequence $Y\in\mathbb{R}^{M}$. Under continuous settings, the SSMs are defined according to Eq.1:
$\begin{cases}\dot{h}(t)&=\bm{A}h(t)+\bm{B}x(t)\\ y(t)&=\bm{C}h(t)+\bm{D}x(t),\end{cases}$ (1)
where $x(t)\in\mathbb{R}^{L}$, $y(t)\in\mathbb{R}^{M}$, $h(t)$, and $\dot{h}(t)\in\mathbb{R}^{N}$ stands for the input, output, hidden state, and derivative of hidden state at timestamp $t$, respectively; $\bm{A}\in\mathbb{R}^{N\times N},\bm{B}\in\mathbb{R}^{N\times L},\bm{C}\in\mathbb{R}^{M\times N}$ and $\bm{D}\in\mathbb{R}^{M\times L}$ are learnable model parameters.
In real-world applications, the input sequences are discrete samplings of continuous sequences. According to  Gu et al. (2022), under discrete settings, by applying the zero-order hold technique to Eq.1, it can be reformulated as follows.
$\begin{cases}h_{k}=\bm{\bar{A}}h_{k-1}+\bm{\bar{B}}x_{k}\\ y_{k}=\bm{C}h_{k}\end{cases},$ (2)
where $\bm{\bar{A}}=\exp(\Delta\bm{A}),\bm{\bar{B}}=(\Delta\bm{A})^{-1}(\exp(\Delta\bm{A})-\bm{I})\cdot(\Delta\bm{B})$ and $\Delta$ is the learnable step size in discrete sampling. We can see from Eq.2 that the hidden state is updated according to the input $x(t)$ and last hidden state $h(t-1)$ while the output is generated by the hidden state $h(t)$ and the input $x(t)$ and in Gu et al. (2020), where it introduces High-order Polynomial Projection Operator (Hippo) to achieve longer sequence modeling.
However, it is worth noticing that $\bm{A},\bm{B},\bm{C},\bm{D}$ in Eq.1 and Eq.2 are time-invariant parameters, i.e., they are data-independent parameters and do not change over time. Therefore the model is not capable of assigning different weights at different positions in the input sequence while receiving new inputs. To address this issue, Gu & Dao (2023) proposed Mamba, in which the parameter matrices $\bm{A},\bm{B},\bm{C},\bm{D}$ are input-dependent, thus enhancing the performance of sequence modeling. To tackle the problem of non-parallelization, Gu & Dao (2023) also introduced selective scan mechanism for effective computing. For further performance and efficiency improvements, Dao & Gu (2024) point out that SSMs can be categorized as a variant of linear attention model. In this work, we follow the same architecture of parallel Mamba Blocks as Dao & Gu (2024) and a RMS-norm Zhang & Sennrich (2019) module is added after the parallel Mamba block. The details of the post-normalization Mamba Block (PNM Block) are illustrated in Fig.3a.
2.2 Diffusion Models
Let $x_{t}$ be a sequence of variables for $t=1,2,\cdots,T$. The diffusion process consists of two processes: 1) The forward process without learnable parameters, which transforms the data distribution into a standard Gaussian distribution by gradually adding noise to the data. 2) The reverse process with learnable parameters, which first samples from the standard Gaussian distribution and then progressively denoises the data to approximate the data distribution. The reverse process of diffusion models a parameterized distribution $p_{\theta}$ defined with the following Markov chain to approximate the real data distribution:
$p_{\theta}(x_{0:T})=p(x_{T})\prod_{t=1}^{T}p_{\theta}(x_{t-1}|x_{t}),$ (3)
where $x_{T}\sim\mathcal{N}(0,I)$ denotes the latent variable sampled from standard Gaussian distribution and
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma_{\theta}(x_{t},t)I),$ (4)
The loss function of DDPM aims at minimizing the difference between the noise in the forward process $\epsilon$ and the parameterized noise $\epsilon_{\theta}$ in the reverse process:
$\mathcal{L}_{d}=\mathbb{E}_{x_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(x_{t},t)\|,$ (5)
where $t$ stands for the diffusion time embedding and $x_{t}$ is calculated in the forward process. Please refer to Appendix 7.1 for more details about the diffusion models.
2.3 Problem Formulation
Definition 1 (Time Series).
A time series can be defined as a tuple, denoted as $\tilde{X}=(X,M,T)$, where $X\in\mathbb{R}^{K\times L}$ is the observation matrix with $K$ observations at a time, which are ordered along $L$ time intervals chronologically; $M\in\mathbb{R}^{K\times L}$ is an indicator matrix that indicates whether the observation at $(i,j)$ in $X$ is missing or not: if the observation at position $(i,j)$ is missing, i.e., $X_{i,j}=\text{NA}$, then $M_{i,j}=1$, otherwise, $M_{i,j}=0$; $T\in\mathbb{R}^{L}$ is the time stamps of the time series.
Definition 2 (Probabilistic Time Series Imputation).
Given an incomplete time series $\tilde{X}=(X,M,T)$, where $\sum M<K\cdot L$, the problem of probabilistic time series imputation is to learn an imputation function $\mathcal{M}_{\theta}$, such that
$\bar{X}=\mathcal{M}_{\theta}(\tilde{X}),$ (6)
where $\bar{X}\in\mathbb{R}^{K\times L}$ is the imputed time series, where $\bar{X}_{i,j}=\mu_{i,j}\pm\sigma_{i,j}$ denotes the probabilistic output if $M_{i,j}=1$, otherwise $\bar{X}_{i,j}={X}_{i,j}$.
3 Methodology
3.1 Diffusion Models for Time Series Imputation
When dealing with time series imputation using diffusion models, consider a time series $\tilde{X}$, our goal is to model the posterior $P(\bar{X}|X,M,T)$. To make the modeled posterior more precisely, it is natural to introduce conditions to introduce the diffusion process. Considering the short range and long range inter-dependencies within time series, maximizing the observed values utilized in the diffusion process can effectively improve the performance of the imputation results. On the other hand, due to the fact that all the observed values are utilized as condition inputs in the diffusion process, we do not apply any extra process to the observed values to avoid the error accumulation caused by information propagation, the observed values $X_{o}^{c}$ are condition inputs for the diffusion process. Thus, the reverse process in Eq.3 is modified to a conditional form with time-series inputs:
$p_{\theta}(X^{m}_{0:T}|X_{0},X_{o}^{c})=p(X_{T}^{m})\prod_{t=1}^{T}p_{\theta}(X_{t-1}^{m}|X_{t}^{m},X_{o}^{c}),$ (7)
where $X_{T}^{m}\sim\mathcal{N}(0,I)$, $X_{t}^{m}$ denotes the sequence of latent variables in the diffusion process and $t\in\{1,2,\cdots,T\}$ is the diffusion time steps. Eq.4 is reformulated as:
$p_{\theta}(X_{t-1}^{m}|X_{t}^{m},X_{o}^{c})=\mathcal{N}(X_{t-1}^{m};\mu_{\theta}(X_{t}^{m},t|X_{o}^{c}),\sigma_{\theta}(X_{t}^{m},t|X_{o}^{c})\bm{I}),$ (8)
the parameterized mean turns to:
$\mu_{\theta}(X_{t},t)=\frac{1}{\alpha_{t}}\left(X_{t}-\frac{\beta_{t}}{\sqrt{1-\alpha_{t}}}\epsilon_{\theta}(X_{t},t|X_{o}^{c})\right),$ (9)
where
$X_{t}=\sqrt{\alpha_{t}}X_{0}+(1-\alpha_{t})\epsilon,$ (10)
and $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is a predefined variance scheduler and $\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{t})$, hence we get the conditional diffusion loss for time series imputation task:
$\mathcal{L}=\mathbb{E}_{X_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(X_{t},t|X_{o}^{c})\|=\mathbb{E}_{X_{0},\epsilon}\|\epsilon-\epsilon_{\theta}(\sqrt{\alpha_{t}}X_{0}+(1-\alpha_{t})\epsilon,t|X_{o}^{c})\|,$ (11)
where $\epsilon\sim\mathcal{N}(0,I)$.
In the real world, the imputation problem encounters various complexities, such as different ratios of missing data, the positions of missing values within the sequence and the distribution of missing data. To simulate various complex missing situations in real-world scenarios, we adopt a self-supervised approach for training, i.e., applying a predefined mask to the complete dataset to construct corresponding dataset with missing data. We follow the same mask strategies in Alcaraz & Strodthoff (2023), including Random Missing (RM) which corresponds to the situation of uniformly random missing values, Random Block Missing (RBM) which corresponds to the situation of continuous missing values (missing intervals) in different channels and Blackout Missing (BM) which contains missing intervals at the same timestamps among different channels.
Figure 1: The self-supervised framework and training process of DiffImp. First, some of observed values are masked following the same missing pattern as the missing values (in red) to get masked targets ($X_{0}$, in magenta) and the condition input ($X_{o}^{c}$, in blue). The noisy input is obtain from $X_{0}$ and $\epsilon$ (in orange) sampled from $\mathcal{N}(0,I)$ The objective of the network is to minimize the difference between the parameterized noise $\epsilon_{\theta}(X_{t},t)$ and $\epsilon$. Solid lines in each time series represent observed values, while dashed lines represent missing values.
3.2 Model Architecture
The Overall Module Architecture Fig.1 illustrates the overall self-supervised framework and training process of our model. We first mask part of the observed values according to the pattern of missing values, where the masked values serve as the imputation target $X_{0}$ during training. The remaining observed values form the conditional input $X_{o}^{c}$ for the noise prediction network $\epsilon_{\theta}$. We then combine $X_{0}$ with noise $\epsilon$ sampled from a standard normal distribution to obtain the noisy input $X_{t}$. Both $X_{o}^{c}$, $X_{t}$, and the diffusion step $t$ are fed into the noise prediction network $\epsilon_{\theta}$ to get the parameterized noise. The network minimizes the difference between $\epsilon_{\theta}$ and $\epsilon$ according to Eq.11.
As shown in Fig.2, the forward process of $\epsilon_{\theta}$ are as follows: For each diffusion step, the input consists of the following parts: noisy input $X_{t}$, the condition input $X_{o}^{c}$ and the diffusion step $t$. To begin with, the inputs are embedded to the latent diffusion space. The embedding module of noisy inputs and condition inputs share a similar model structure, which consists of a linear projection module followed by an SMM block in Fig.3b. The SMM block is composed of stacks of Bidirectional Attention Mamba (BAM) blocks and Channel Mamba Blocks (CMB), which is introduced in the next part. Due to the relatively limited information from $t$, the embedding module of $t$ only consists of linear projection modules. After the embedding step, the embedded diffusion step is concatenated with the input embeddings. The concatenated embeddings are fed in to a SMM module. Then the output of the SMM module is concatenated with the condition embeddings. After feeding the final embeddings to another SMM module and final projection module, we can get the noise predictions $\epsilon_{\theta}(X_{t},t)$. The training and sampling algorithm is detailed in Alg.1 and Alg.2.
Figure 2: Architecture of $\epsilon_{\theta}$ in DiffImp
Mamba Encoders for Bidirectional Modeling For probabilistic time series imputation tasks, the objective is to attain a more precise posterior estimation for the missing points contingent upon the observed points. Therefore, our proposed module should achieve two key objectives: Firstly, it should possess bidirectional analysis capability, which means that the model should be able to capture dependencies in both the forward and reverse temporal directions. Secondly, considering that the known points at different positions relative to the missing point have varying distances, the model should assign different weights to difference timestamps. To address these issues, we devise a bidirectional attention Mamba module (BAM). BAM takes the representations from previous layers as input, which are then fed into two distinct PNM modules (Fig.3a), enabling the model to capture bidirectional dependencies. More specifically, temporal attention is implemented by assigning different values to various time steps in the sequence, where the temporal attention module also receives the previous layer’s representation and learns the weights for different timestamps. The details of BAM are shown in Fig.3d.
Mamba Encoders for Inter-channel Modeling In the context of multivariate time series, inter-dependencies exist among variables across different channels. The effective modeling of these inter-channel correlations is instrumental in capturing the intrinsic characteristics of the time series more adeptly. Additionally, when analyzing the relationships between channels, the order of the channels does not exhibit the sequential dependencies as that among timestamps. Consequently, we employ unidirectional channel dependency modelling architecture, termed as Channel Mamba Block (CMB). We first transpose the input time series representation for processing on the channel dimension. The transposed representations are then subjected to a normalization module and processed through a PNM block, yielding a more profound feature representation. The details of CMB are presented in Fig.3c.
Figure 3: Details of PNM, SMM, CMB, BAM block in the noise prediction module. (a) PNM: backbone module based on Mamba. (b) SMM: core components of noise prediction module, composed of stacks of BAM and CMB. (c) CMB: unidirectional module for inter-channel dependency modeling. (d) BAM: bidirectional module with temporal attention for intra-channel, multi-range dependency modeling.
Algorithm 1 Training Procedure of DiffImp
1:  Input: Observed sequence $x_{0}$, number of iterations $N$, variance scheduler $\beta_{t}$
2:  Output: Denoising function $\epsilon_{\theta}$
3:  For i = 1 to $N$ do:
4:   tsimilar-$t\sim$ Uniform($(\{1,2,\cdots,T\})$
5:   $\epsilon\sim\mathcal{N}(0,I)$
6:   Calculate diffusion targets $x_{t}$ according to Eq.10
7:   Take gradient step on
$\nabla_{\theta}(\|\epsilon-\epsilon_{\theta}(x_{t},t|X_{0})\|)$
according to Eq.11
8:  End For
Algorithm 2 Sampling Procedure of DiffImp
1:  Input: Trained denoising function $\epsilon_{\theta}$, sampling step $T$
2:  Output: Mean prediction $x_{0}$
3:  For $t=T,T-1,\cdots,1$ do:
4:   $z\sim\mathcal{N}(0,I)$ if $t>1$ else $z=0$
5:   $x_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\right)\epsilon_{\theta}(x_{t},t)+\sigma_{t}z$
6:  End For
Complexity Analysis While dealing with the input sequences, the core component of our module is the PNM module in Fig.3 and the self-attention module in the Transformer architecture, respectively. In this part, we will give a brief analysis about the time and space complexity in the SSM module and self-attention module [^1] We do not take the time and space complexity of MLPs before the self-attention module or SSM module into consideration.. The time complexity of self-attention module is $O(CL^{2})$ and the space complexity is $O(L^{2}+CL)$, where $L$ is the length of the input sequence and $C$ is the channel of the input sequence.
In our method, the forward process described in Eq.2 is implemented by converting the process to multiplications of structured matrices, which is of time complexity $O(NCL)$ and of space complexity $O(CL+N(C+L))$ ($N$ is a constant number and set as $16$ by default). This indicates that our model is of linear time and space complexity with respect to the sequence length $L$, which ensures scalability and reduces memory cost for longer sequences.
4 Experiments
4.1 Experiment Settings
Datasets and Experimental Settings We conduct experiments on three real-world datasets to validate the effectiveness of our approach. These datasets span multiple domains, namely Electricity dataset Asuncion & Newman (2007), MuJoCo dataset Rubanova et al. (2019b) and ETTm1 dataset Zhou et al. (2021).
All experiments are conducted using PyTorch Paszke et al. (2019) in Python 3.9 and execute on an NVIDIA RTX3090 GPU. The training process is guided by Eq.11, employing the ADAM optimizer Kingma & Ba (2015) with a learning rate of $2\times 10^{-4}$. More details about the datasets and experimental settings can be found in the Appendix.
Evaluation Metrics and Baselines To achieve an extensive evaluation of imputation performance, diverse metrics are utilized for evaluating deterministic imputation results, namely Mean Absolute Error (MAE), Mean Squared Error (MSE), and Root Mean Square Error (RMSE). Due to reproducibility reasons of baselines, we compare with different baselines and report different metrics for different datasets. The datasets and corresponding baseline and metrics are listed in Table.2. We follow the same settings and dataset preprocessings as Alcaraz & Strodthoff (2023) and collect all the baseline results from the same paper.
As for the evaluation of probabilistic imputation, we calculate the Continuous Ranked Probabilistic Score-sum (CRPS-sum) on the electricity dataset. The CRPS-sum results are collected from Yan et al. (2024). In all the tables of our experiment results, the best results are in bold and second best results are underlined. All the deterministic metrics are maintained by running the experiment for 3 times and CRPS-sum is obtained by 10 runs.
Table 2: Datasets and corresponding evaluation metrics and baselines for time series imputation and forecasting task.
Dataset Task Metric Baseline
Electricity Imputation MAE; RMSE; MRE
M-RNN Yoon et al. (2019); GP-VAE Fortuin et al. (2020);
BRITS Cao et al. (2018); SAITS Du et al. (2023); CSDI; SSSD
MuJoCo Imputation MSE
RNN GRU-D Che et al. (2018); ODE-RNN Rubanova et al. (2019a);
NeuralCDE Morrill et al. (2021); Latent-ODE Rubanova et al. (2019a);
NAOMI Liu et al. (2019); NRTSI Shan et al. (2023a); CSDI; SSSD
ETTm1 Forecasting MAE; MSE
LSTNet Lai et al. (2018); LSTM Bahdanau et al. (2015);
Reformer Kitaev et al. (2020); LogTrans Li et al. (2019);
Informer Zhou et al. (2021); Autoformer Wu et al. (2021); CSDI; SSSD
4.2 Time Series Imputation
[…]
RNN GRU-D 1.134e-2 1.421e-2 1.968e-2
ODE-RNN 9.86e-3 1.209e-2 1.647e-2
NeuralCDE 8.35e-3 1.071e-2 1.352e-2
Latent-ODE 3.00e-3 2.95e-3 3.60e-3
NAOMI 1.46e-3 2.32e-3 4.42e-3
NRTSI 6.3e-4 1.22e-3 4.06e-3
CSDI 2.4e-4$\pm$3e-5 6.1e-4$\pm$1.0e-4 4.84e-3$\pm$2e-5
SSSD 5.9e-4$\pm$8e-5 1e-3$\pm$5e-5 1.90e-3$\pm$3e-5
DiffImp (Ours) 2.7e-4$\pm$1e-5 3.16e-4$\pm$9.77e-6 6.5e-4$\pm$1e-4
Table.4 shows the experimental results on the Electricity dataset, where we apply the RM missing pattern with missing ratios of 10%, 30%, and 50%. We achieve the best results across all metrics with a 30% missing ratio, significantly outperforming other methods. In the experiments with 10% and 50% missing ratios, we obtain results with only a slight gap to the SOTA models.
Table 4: MAE and RMSE results on Electricity Dataset
[…]
Model MAE RMSE MRE MAE RMSE MRE MAE RMSE MRE
M-RNN 1.244 1.867 66.6% 1.258 1.876 67.3% 1.283 1.902 68.7%
GP-VAE 1.094 1.565 58.6% 1.057 1.571 56.6% 1.097 1.572 58.8%
BRITS 0.847 1.322 45.3% 0.943 1.435 50.4% 1.037 1.538 55.5%
SAITS 0.735 1.162 39.4% 0.790 1.223 42.3% 0.876 1.377 46.9%
CSDI 1.510$\pm$3e-3 15.012$\pm$4e-2 81.10$\pm$1e-10.921$\pm$8e-3 8.732$\pm$7e-2 49.27$\pm$4e-10.278$\pm$4e-3 2.371$\pm$3e-2 14.93$\pm$1e-1%
SSSD 0.345$\pm$1e-4 0.554$\pm$5e-5 18.4$\pm$5e-30.407$\pm$5e-4 0.625$\pm$1e-4 21.8$\pm$00.532$\pm$1e-4 0.821$\pm$1e-4 28.5$\pm$1e-2%
DiffImp (Ours) 0.378$\pm$6e-4 0.522$\pm$3e-3 20.2$\pm$1e-20.348$\pm$1e-3 0.496$\pm$2e-3 18.6$\pm$1e-10.546$\pm$3e-3 0.837$\pm$7e-3 29.2$\pm$2e-1%
Probabilistic Imputation Results Table.5 presents a comparison of our method with other probabilistic time series imputation methods based on the CRPS-sum metric. The baselines for CRPS-sum include Tashiro et al. (2021); Chen et al. (2023); Alcaraz & Strodthoff (2023); Yan et al. (2024); Kollovieh et al. (2023). The experimental results show that our method achieves a 21.4% performance improvement compared to the second-best method. This indicates that our method models the data distribution of the sequence more accurately than other baseline methods.
Table 5: The CRPS-sum results on electricity dataset
Model CSDI CSBI SSSD TS-Diff D3M DiffImp(Ours)
CRPS-sum 2.14e-2$\pm$8e-3 2.19e-2$\pm$7e-3 1.96e-2$\pm$1e-3 2.23e-2$\pm$6e-3 1.92e-2$\pm$4e-3 1.51e-2$\pm$4e-4
4.3 Time Series Forecasting
As mentioned in 3.1, the probabilistic time series forecasting problem can be treated as a variant of the probabilistic time series imputation problem (as a special case of the missing manner BM). Therefore, we also conduct experiments to validate the effectiveness of our experiments on probabilistic time series forecasting task. Following the setup in previous works, we test five different forecasting horizons: 24, 48, 96, 288, and 672 time steps, with corresponding conditional lengths (i.e., the length of observed sequence) of 96, 48, 284, 288, and 384 time steps.
Table.6 presents the experimental results on the ETTm1 dataset. Our method achieves state-of-the-art performance on prediction length of 24 and 96, outperforms other imputation-based algorithms at the prediction length of 672, and shows only a slight gap compared to the best imputation-based algorithms at the prediction length of 48 and 288.
Table 6: MSE and MAE results on ETTm1 dataset
Forecasting Length 24 48 96 288 672
Model MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE
LSTNet 1.170 1.968 1.215 1.999 1.542 2.762 2.076 1.257 2.941 1.917
LSTMa 0.629 0.621 0.939 1.392 0.913 1.339 1.124 1.740 1.555 2.736
Reformer 0.607 0.724 0.777 1.098 0.945 1.433 1.094 1.820 1.232 2.187
LogTrans 0.412 0.419 0.583 0.507 0.792 0.768 1.320 1.462 1.461 1.669
Informer 0.369 0.323 0.503 0.494 0.614 0.678 0.786 1.056 0.926 1.192
CSDI 0.370$\pm$3e-3 0.354$\pm$1.5e-2 0.546$\pm$2e-3 0.750$\pm$4e-3 0.756$\pm$1.1e-2 1.468$\pm$4.7e-2 0.530$\pm$4e-3 0.608$\pm$3.5e-2 0.891$\pm$3.7e-2 0.946$\pm$5.1e-2
Autoformer 0.403 0.383 0.453 0.454 0.463 0.481 0.528 0.634 0.542 0.606
SSSD 0.361$\pm$6e-3 0.351$\pm$9e-3 0.479$\pm$8e-3 0.612$\pm$2e-3 0.547$\pm$1.2e-2 0.538$\pm$1.3e-2 0.648$\pm$1.0e-2 0.797$\pm$5e-3 0.783$\pm$6.6e-2 0.804$\pm$4.5e-2
DiffImp (Ours) 0.282$\pm$1.8e-2 0.331$\pm$9.9e-3 0.679$\pm$5.6e-3 0.548$\pm$5.6e-4 0.3906$\pm$1.3e-2 0.4211$\pm$8.5e-3 0.621$\pm$2.1e-3 0.741$\pm$3.3e-3 0.683$\pm$3.1e-3 0.783$\pm$6.8e-3
4.4 Visualization Results
Figure 4: Visualized results of probabilistic time series imputation on MuJoCo dataset.
Fig.4 shows the visualization results for channel 5 and channel 7 on the MuJoCo dataset with a 90% missing ratio. From the figure, we can see that almost all ground truth values for the points to be imputed fall within the 95% confidence interval, and most of the ground truth values are within the 50% confidence interval, which demonstrates the effectiveness of our method.
4.5 Sampling Time Analysis
Table presents a comparison of sampling time between our method and other backbone-based methods across different datasets. We find that, in terms of sampling time, our method performs better compared to transformer-based method and previous SSM-based method, demonstrating the efficiency of our backbone choice. The data in the table represent the time of a single sample process.
Table 7: Sampling time of CSDI, SSSD and DiffImp on Electricity and MuJoCo dataset, - denotes the corresponding code is note provided in original papers.
CSDI SSSD DiffImp
Electricity 1.43s - 1.012s
MuJoCo - 4.16s 1.01s
4.6 Ablation Studies
To validate the effectiveness of the proposed module, we conduct ablation experiments on the following aspects: 1) the bidirectional modeling 2) the temporal attention mechanism 3) the inter-channel multivariate dependencies. All experiments are conducted on the MuJoCo dataset with the missing ratio 90%. During ablation experiments, we find out that our model converges much slower than other models in the ablation experiment, so we train till all models are converged (for same number of iterations, even if it has already been converged). The hyperparameters in the ablation studies are presented in the appendix.
The results are shown in Table.8. It can be observed that the module equipped with BAM and CMB block performs the best, significantly outperforming the results of removing any one of these components across all four metrics. The temporal attention module has the largest impact on the model, and its removal leads to a significant performance drop. Similarly, removing the CMB module also results in a notable degradation in performance. On the other hand, adjusting the BAM module to its unidirectional form also causes some degree of performance decrease. This fully demonstrates the effectiveness of our proposed blocks.
Table 8: Experimental results of Ablation Study
Time Modeling Temporal Attention Inter-Channel Dependency MSE MAE MRE RMSE
Bidirectional Yes Yes 5.46e-4$\pm$1.6e-5 1.17e-2$\pm$7.4e-5 1.21$\pm$7.5e-32.33e-2$\pm$3.1e-4
Forward Yes Yes 7.19e-4$\pm$2.0e-5 1.26e-2$\pm$2.1e-4 1.29$\pm$2.2e-22.67e-2$\pm$2.9e-4
Forward Yes No 7.48e-4$\pm$9.5e-5 1.23e-2$\pm$3.5e-4 1.23$\pm$3.5e-22.71e-2$\pm$1.5e-3
Backward Yes Yes 7.24e-4$\pm$7.3e-5 1.30e-2$\pm$4.2e-4 1.30$\pm$4.2e-22.69e-2$\pm$1.3e-3
Backward Yes No 8.39e-4$\pm$6.1e-5 1.46e-2$\pm$3.8e-4 1.46$\pm$3.8e-22.89e-2$\pm$1.0e-3
Bidirectional Yes No 8.85e-4$\pm$2.8e-5 1.40e-2$\pm$3.1e-4 1.44$\pm$3.4e-22.97e-2$\pm$4.8e-4
Bidirectional No Yes 9.66e-4$\pm$9.5e-5 1.53e-2$\pm$3.3e-4 1.57$\pm$3.5e-23.09e-2$\pm$1.3e-3
5 Conclusion and Future Work
In this paper, we propose DiffImp, a time series imputation model based on DDPM and Mamba backbone, which incorporates bidirectional information flow, temporal attention and inter-variable dependencies. DiffImp enables efficient time series modeling with linear complexity. Experimental results demonstrate that DiffImp achieves superior performance across multiple datasets, various missing patterns, and different missing ratios.
For future work, one possible direction is to further reduce the time complexity of the sampling process while already lowering the complexity of time series modeling, in order to enhance the model’s inference efficiency. Another possible direction is to extend the application of diffusion models by applying DiffImp to other time series downstream tasks and time series representation learning tasks.
6 Reproducibility
To ensure reproducibility and facilitate experimentation, datasets and code are available at: https://anonymous.4open.science/r/DiffImp-843F.